OpenStudy (anonymous):

solve this differential equation!?! (x^2 + a^2)y' = xy please help..

OpenStudy (anonymous):

dy/y = x/(x^2+a^2) dx integral each side...

OpenStudy (watchmath):

\(\int\frac{dy}{y}=\int\frac{x\,dx}{a^2+x^2}\) \(\ln|y|=\frac{1}{2}\ln(a^2+x^2)+C\)

OpenStudy (anonymous):

oo ok..and we just solve for y?

OpenStudy (anonymous):

use e

OpenStudy (watchmath):

\(|y|=Ae^{\ln(\sqrt{a^2+x^2})}=A\sqrt{a^2+x^2}\) \(y=\pm A\sqrt{a^2+x^2}=A\sqrt{a^2+x^2}\)

OpenStudy (anonymous):

what is the capital A??

OpenStudy (watchmath):

That is to replace \(e^C\) which is just another constant.

OpenStudy (anonymous):

oo for e^(1/2)..?

OpenStudy (anonymous):

no for e^c

OpenStudy (watchmath):

\(\frac{1}{2}\ln (a^2+x^2)=\ln(a^2+x^2)^{1/2}\) by the property of \(\ln\).

OpenStudy (anonymous):

yes..i understand it now thanks guys!

OpenStudy (watchmath):

I think we need to add small detail that \(A >0\).